OpenStudy (anonymous):

Evaluate the function rule y = 6 x 4x for x = -3 A: 3/32 B:3/8 C: -72 D:3/128 WILL MEDAL AND FAN

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

\[y = 6 \times (4x)\]

OpenStudy (anonymous):

Put \(x=-3\), \[y = 6 \times (4 \times \color{red}{(-3)}) \\ y = 6 \times (-12) \\ y = ??\]

OpenStudy (anonymous):

Find out the answer now.. What is \(6 \times (-12) = ??\)

OpenStudy (anonymous):

-72 but the thing said I was incorrect when I said that @eta

OpenStudy (anonymous):

What? C is not right??

OpenStudy (anonymous):

Is this your equation?? \[y = 6 \times (4x)\]

OpenStudy (anonymous):

No I was pretty sure it was but it said incorrect the first time I said that

OpenStudy (anonymous):

Can you go for it once again?

OpenStudy (anonymous):

\[y = 6 \times 4x for x = -3\]

OpenStudy (anonymous):

@eta

OpenStudy (anonymous):

Then we are right according to if that is your question and we can't do more in this.. :(

OpenStudy (anonymous):

Well we can move on to another question?